# The dark timbre of gravitational waves

### v1

@@ Abstract
Detectability and prospects – To detect this effect, we require that the signal with the shifted timbre can not fit well with a template using the standard timbre. In general, the wellness of a template fit is given by Jaranowski and Krolak (2005)
@@
where $A$ characterizes the amplitude of the signal and ${\rm SNR}_{2}$ is the signal-to-noise ratio of the second harmonic. The template is parameterized by an amplitude $\lambda\approx 1$, an eccentricity $e^{\prime}\approx e$, and a phase $\phi\approx 0$. The function $g_{n}(e)$ is the relative power radiated in the $n$th harmonic by a binary with eccentricity $e$ Peters and Mathews (1963). We show $g_{n}(e)$ for some of the lowest harmonics in the lower panel of Fig. 3.
For the spectral sensitivity $S(f)$ we use the proposed spectral sensitivity for LISA Amaro-Seoane et al. (2017); Robson et al. (2019) including the galactic and extragalactic foregrounds Lewicki and Vaskonen (2023). We considered sources at $z_{s}=0.5$ and at $z_{s}=1$ with $f_{2}=2\,$mHz. To assess the detectability of the shifted timbre effect, we consider only the harmonics $n=2-7$ which dominate the signal for $e<0.6$, as seen in the lower panel of Fig. 3. As benchmarks, we consider 1) the mean values for $F_{i,2}$ and 2) the best out of 100 events. We determine how large ${\rm SNR_{2}}$ needs to be for a given $e$ so that the shifted timbre is detectable. We use the threshold $\Delta\chi^{2}>8.02$ corresponding to $2\sigma$ CL for a template with three parameters. In the upper panel of Fig. 3, we show that if we consider $z_{s}=0.5$ (full lines) then ${\rm SNR_{2}}\sim 10^{4}$ is required if $e=0.3$ but for $e=0.5$ it is enough to have ${\rm SNR_{2}}\lesssim 10^{3}$. The required ${\rm SNR_{2}}$ is about a factor 3 smaller for the best out of 100 events than for the mean. These SNRs seem within the reach of LISA for sufficiently nearby sources.
The precision at which the amplitudes of a harmonic $i$ can be resolved is
$\frac{\sigma_{A_{i}}}{A_{i}}=\frac{1}{{\rm SNR}_{i}}\,,$ (15)
@@
which is easy to show from Fisher analysis Poisson and Will (1995). This implies that a higher ${\rm SNR}_{2}$ is necessary to resolve the relative amplification factors $F_{i,j}$ than to detect the dark timbre. Optimistically, if a population of eccentric binaries is detected with ${\rm SNR}_{i}\sim 10^{4}$ it would be possible to reconstruct the PDFs of $F_{i,j}$ and probe the HMF down to $M_{\rm v}\ll 10\,M_{\odot}$. However, already the detection of one sufficiently eccentric binary with ${\rm SNR}_{2}\gtrsim 10^{3}$ would tell about the existence of halos with $M_{\rm v}\lesssim 10\,M_{\odot}$.
@@
Conclusions – We have estimated for the first time the integrated lensing effect of light halos on the mid-frequency (mHz) gravitational wave signals. We have proposed to test these effects using the different harmonics of an eccentric binary. We have shown that the wave optics effects due to the low mass dark matter halos, $M_{\rm v}\lesssim 10\,M_{\odot}$, induce frequency-dependent changes in the amplitude and phase of the harmonics, the timbre of the signal. This shifted timbre is detectable in the 7 dominant harmonics of the signal for signal-to-noise ratios between $500-10^{4}$ if the binary eccentricity is $e=0.3-0.6$. If such binaries exist, LISA could probe the shifted timbre. This would open a new avenue to test low-mass dark matter halos and thus provide new insights into the nature of dark matter.

### v2

@@ IV Detectability and prospects
To detect the dark timbre effect, we require that the signal with the shifted timbre cannot be fitted well with a template using the standard timbre. In general, the wellness of a template fit is given by Jaranowski and Krolak (2005)
@@
where $A$ characterizes the amplitude of the signal and ${\rm SNR}_{2}$ is the signal-to-noise ratio of the second harmonic. The template is parameterized by an amplitude $\lambda\approx 1$, an eccentricity $e^{\prime}\approx e$, and a phase $\phi\approx 0$. The function $g_{n}(e)$ gives the relative power radiated in the $n$th harmonic by a binary with eccentricity $e$ Peters and Mathews (1963). We show $g_{n}(e)$ for some of the lowest harmonics in the lower panel of Fig. 3.
We use the expected LISA spectral sensitivity $S(f)$ Amaro-Seoane et al. (2017); Robson et al. (2019) and include noise arising from the galactic and extragalactic foregrounds Lewicki and Vaskonen (2023). We considered sources at $z_{s}=0.5$ and at $z_{s}=1$ with $f_{2}=2\,$mHz. To assess the detectability of the shifted timbre effect, we consider only the harmonics $n=2-7$ that dominate the signal for $e\lesssim 0.6$, as seen in the lower panel of Fig. 3. We determine how much ${\rm SNR_{2}}$ and $e$ are required to detect the shifted timbre. We use the threshold $\Delta\chi^{2}>8.02$ corresponding to $2\sigma$ CL (Confidence Level) for a template with three parameters. The results are shown in the upper panel of Fig. 3. For example, for $z_{s}=0.5$ and $e=0.3$ the dark timbre is detected if ${\rm SNR_{2}}\gtrsim 3000$ while for $e=0.5$ it is enough to have ${\rm SNR_{2}}\lesssim 500$. For $z_{s}=1$, approximately factor four smaller ${\rm SNR_{2}}$ is sufficient. An example of a source with approximately these properties could be an IMBH binary with $\mathcal{M}=\mathcal{O}(10^{3}/M_{\odot})$ and $\eta\sim 0.1$. In scenarios with light supermassive BH seeds, the number of detectable binaries of this type would be $\mathcal{O}(100)/{\rm year}$ and they would enter the LISA band at $\mathcal{O}({\rm years})$ before merger Ellis et al. (2024). For some time, they would be approximately monochromatic with SNR of $\mathcal{O}(10^{3})$. As shown in Porter and Sesana (2010), also the referred eccentricities are realistic.
The precision at which the amplitudes of a harmonic $i$ can be resolved is approximately
$\frac{\sigma_{A_{i}}}{A_{i}}\sim\frac{1}{{\rm SNR}_{i}}\,,$ (15)
@@
which is easy to show from Fisher analysis Poisson and Will (1995). This implies that a higher ${\rm SNR}_{2}$ is necessary to resolve the relative amplification factors $F_{i,j}$ than to detect the dark timbre. Optimistically, if a population of eccentric binaries is detected with ${\rm SNR}_{i}\sim 10^{3}$ it would be possible to reconstruct the PDFs (Probability Density Functions) of $F_{i,j}$ and probe the HMF down to $M_{\rm v}\ll 10\,M_{\odot}$. However, already the detection of one sufficiently eccentric binary with ${\rm SNR}_{2}\gtrsim 10^{2}$ at $z_{s}=0.5$ would tell about the existence of halos with $M_{\rm v}\lesssim 10\,M_{\odot}$.
We note that since the dark timbre effect is very small, corrections beyond the Newtonian order Peters and Mathews (1963) in the different harmonics coefficients may become relevant, although we assume that the binaries are very far from merging, furthermore considering more realistic waveforms with higher modes would potentially help in the detectability Çalışkan et al. (2023a). Moreover, detecting the dark timbre requires identifying the frequency harmonics. LISA is expected to observe several nearly monochromatic GWs simultaneously at different frequencies. Some of these GWs may be harmonics from an eccentric IMBH binary, while others may originate, e.g., from galactic WD binaries Korol et al. (2017). Consequently, resolving which of these GWs are harmonics emitted by the same source can be challenging. The effects of higher modes and the identification problem of the frequency harmonics are left for future studies.
V Conclusions
@@
We have estimated for the first time the integrated lensing effect of light halos on the mid-frequency (mHz) GW signals. We have proposed to test these effects using the different harmonics of an eccentric binary. We have shown that the wave optics effects due to the low-mass DM halos, $M_{\rm v}\lesssim 10\,M_{\odot}$, induce frequency-dependent changes in the amplitude and phase of the harmonics – the timbre of the signal. This shifted timbre is detectable in the 7 dominant harmonics of the signal for signal-to-noise ratios between $100-10^{3}$ if the binary eccentricity is $e=0.3-0.6$ and it is $z_{s}=0.5$. If such binaries exist, LISA could probe the dark timbre. This would open a new avenue to test low-mass DM halos and thus provide new insights into the nature of DM.
@@ Acknowledgements.
Appendix A Validity of the Born and eikonal approximation
In the following, we make explicit the assumptions behind the formula of the amplification factor $F$ used in the main text and showcase their validity.
To compute the scatter of a GW from the DM halo we will consider the metric around the halo to be given by
${\rm d}s^{2}=-(1+2U(\boldsymbol{x})){\rm d}t^{2}+a^{2}(1-2U(\boldsymbol{x})){\rm d}\boldsymbol{x}^{2}\,,$ (16)
where $U$ is the gravitational potential of the halo and $a$ denotes the scale factor that we consider as a constant fixed by the halo redshift. Under the Born approximation, we consider only the leading order terms in $U\ll 1$. Moreover, we neglect the change in the polarization of the GW which is of the order $U\ll 1$ and consider only its amplitude $\phi(t,\boldsymbol{x})$. Then, the Fourier transform $\tilde{\phi}(\omega,r)$ of the scalar wave follows the wave equation
$\left(a^{-2}\nabla^{2}+\omega^{2}\right)\tilde{\phi}(\omega,\boldsymbol{x})=4\omega^{2}\,U\tilde{\phi}(\omega,\boldsymbol{x})\,.$ (17)
We want to express the solution as
$\tilde{\phi}(\omega,\boldsymbol{x})=F(\omega,\boldsymbol{x})\tilde{\phi}_{\rm NL}(\omega,\boldsymbol{x})\,,$ (18)
where the $\tilde{\phi}_{\rm NL}(\omega,\boldsymbol{x})$ is the solution of Eq. (17) for $U=0$. Plugging (18) into the equation of motion (17) we get that the amplification function $F(\omega,r)$ has to fulfill
$\frac{1}{(a\,r)^{2}}\nabla^{2}_{\theta}F+\frac{1}{a^{2}}\frac{\partial^{2}F}{\partial\,r^{2}}+\frac{2i\omega}{a}\frac{\partial F}{\partial r}=4\omega^{2}UF\,.$ (19)
Assuming that
$\left|\frac{1}{a^{2}}\frac{\partial^{2}F}{\partial\,r^{2}}\right|\ll\left|\frac{2i\omega}{a}\frac{\partial F}{\partial r}\right|\,,$ (20)
Eq. (19) simplifies to
$\frac{i}{a}\frac{\partial F}{\partial r}=\frac{1}{2\omega(a\,r)^{2}}\nabla^{2}_{\theta}F+2\omega UF\,.$ (21)
which is a Schrödinger equation that can be solved with the path integral approach that gives the formula of $F$ we use throughout the paper.
In the lower panel of Fig. 4 we plot the NFW gravitational potential,
$U(r)=-\frac{4\pi\rho_{0}r_{s}^{3}}{r}\log{\left(1+\frac{r}{r_{s}}\right)}\,,$ (22)
to show that the Born approximation is valid for the impact parameters we typically consider for the dark timbre. In the upper panel we demonstrate the validity of the eikonal approximation.
Figure 4: Upper panel: Comparison of the second and first derivatives of $F$. Lower panel: Values of $U$ for the typical impact parameters that contribute to the dark timbre.
Appendix B Multilens computation
Figure 5: Comparison between the amplification function for a single lens computed without (solid contours) and with (dashed contours) the expansion in $\omega\psi_{j}$.
The solution of Eq. (21) is given by the path integral
$F=\int{\mathcal{D}}\boldsymbol{\theta}\,e^{i\int_{0}^{r_{N+1}}{\rm d}r\,\mathcal{L}}\,,$ (23)
where $\boldsymbol{\theta}=\theta(\cos{(\phi)},\sin{(\phi)})$ parametrizes the paths that the GW can take from the source at $r=0$, $\boldsymbol{\theta}=0$ to the observer at $r=r_{N+1}$, $\boldsymbol{\theta}=\boldsymbol{\theta}_{N+1}=0$ and
$\mathcal{L}=\frac{1}{2}\omega(ar)^{2}|\dot{\boldsymbol{\theta}}(r)|^{2}-2\omega U\,.$ (24)
We divide the path into $N$ slices perpendicular to the radial vector from the source to the detector at distances $r=r_{j}$ of thickness $\epsilon$, so that $r_{N+1}=(N+1)\epsilon$. Some of these slices include a lens. In the thin lens approximation, we project the gravitational potentials of each lens to a plane,
$U(r,\boldsymbol{\theta})=\frac{1}{2}\sum_{j=1}^{N}\psi_{j}(\boldsymbol{\theta})\delta(r-r_{j})$ (25)
and consider paths that change direction only at the lens planes,
$|\dot{\boldsymbol{\theta}}(r)|=\sum_{j=1}^{N}\Theta(r-r_{j})\Theta(r_{j+1}-r)\frac{r_{j}r_{j+1}}{ar\epsilon}|\boldsymbol{\theta}_{j+1}-\boldsymbol{\theta}_{j}|\,,$ (26)
where $\boldsymbol{\theta}_{j}$ denotes the position where the path crosses the $j$th lens plane and $r_{i,j}\equiv r_{i}r_{j}/(r_{i}-r_{j})$. Then the integrated Lagrangian can be expressed as
$\int_{0}^{r_{N+1}}{\rm d}r\,\mathcal{L}=\omega\sum_{j=1}^{N}\left[\epsilon\frac{r_{j}r_{j+1}}{2}\left|\frac{\boldsymbol{\theta}_{j}-\boldsymbol{\theta}_{j+1}}{\epsilon}\right|^{2}-\Psi_{j}(\boldsymbol{\theta}_{j})\right]\,.$ (27)
and amplification function as
$F=\!\int$ $\!\prod_{j=1}^{N}\frac{\omega r_{j}r_{j+1}}{2\pi i\epsilon}{\rm d}^{2}\boldsymbol{\theta}_{j}$ (28)
×$\times\exp\!\bigg{[}i\omega\bigg{(}\frac{r_{j}r_{j+1}}{2\epsilon}|\boldsymbol{\theta}_{j+1}-\boldsymbol{\theta}_{j}|^{2}-\psi_{j}(\boldsymbol{\theta}_{j})\bigg{)}\bigg{]}.$
Here $\psi_{j}=0$ for those slices that don’t include a lens. Assuming that for all relevant paths $\omega\psi_{j}\ll 1$, we can expand to the lowest order in powers of $\omega\psi_{j}$:
$F\approx$ $\int\!\prod_{j=1}^{N}\frac{\omega r_{j}r_{j+1}}{2\pi i\epsilon}{\rm d}^{2}\boldsymbol{\theta}_{j}\exp\!\bigg{[}i\omega\frac{r_{j}r_{j+1}}{2\epsilon}|\boldsymbol{\theta}_{j+1}-\boldsymbol{\theta}_{j}|^{2}\bigg{]}$ (29)
×$\qquad\times\left[1-i\omega\psi_{j}(\boldsymbol{\theta}_{j})\right]$
$\approx$ $1-i\omega\sum_{l=1}^{N}\int\frac{\omega r_{l}r_{l+1}}{2\pi i\epsilon}{\rm d}^{2}\boldsymbol{\theta}_{l}\left[\prod_{j\neq l}\frac{\omega r_{j}r_{j+1}}{2\pi i\epsilon}{\rm d}^{2}\boldsymbol{\theta}_{j}\right]$
×$\qquad\times\psi_{l}(\boldsymbol{\theta}_{l})\exp\!\bigg{[}i\omega\sum_{j=1}^{N}\frac{r_{j}r_{j+1}}{2\epsilon}|\boldsymbol{\theta}_{j+1}-\boldsymbol{\theta}_{j}|^{2}\bigg{]}.$
The integrals for $j<l$ give unity and Nakamura and Deguchi (1999)
$\sum_{j=l}^{N}\frac{r_{j}r_{j+1}}{2\epsilon}|\boldsymbol{\theta}_{j+1}-\boldsymbol{\theta}_{j}|^{2}=\epsilon\frac{r_{l}r_{N+1}}{r_{l,N+1}}|\boldsymbol{\theta}_{N+1}-\boldsymbol{\theta}_{l}|^{2}$ (30)
$\qquad\qquad\qquad+\sum_{j=l+1}^{N}r_{j}^{2}\frac{r_{l}r_{j+1}}{r_{l,j}}|\boldsymbol{\theta}_{j}-\boldsymbol{u}_{l,j}|^{2}\,,$
where $r_{l,j}=r_{l}-r_{j}$ and $\boldsymbol{u}_{l,j}=[r_{l}\boldsymbol{\theta}_{j}+(j-l)r_{j+1}\boldsymbol{\theta}_{j+1}]/(jr_{l,j+1})$. So, all integrals except that over $\boldsymbol{\theta}_{l}$ are trivial and give $\epsilon r_{N+1}/(r_{l+1}r_{l,N+1})$. This results in
$F$ $\approx 1-i\omega\sum_{l=1}^{N}\int\frac{\omega r_{l}r_{N+1}}{2\pi ir_{l,N+1}}{\rm d}^{2}\boldsymbol{\theta}_{l}\psi_{l}(\boldsymbol{\theta}_{l})$ (31)
×$\qquad\qquad\qquad\times\exp\!\bigg{[}i\omega\frac{r_{l}r_{N+1}}{r_{l,N+1}}|\boldsymbol{\theta}_{N+1}-\boldsymbol{\theta}_{l}|^{2}\bigg{]}$
$\approx 1-\sum_{l=1}^{N}(F_{j}-1)\,.$
In Fig. 5, we show that the expansion in $\omega\psi_{j}\ll 1$ works well for the NFW lenses. The solid and dashed contours are computed without and with the expansion and their excellent agreement is evident.